# Bijective rigidity of uniform Roe algebras and injectivity of the comparison map

### v1

@@ 1.2. Uniform Roe algebras and rigidity
The rigidity problem for uniform Roe algebras (Conjecture 0.1) asks whether Morita equivalences (respectively, isomorphisms) of uniform Roe algebras induce coarse equivalences (respectively, bijective coarse equivalences) between the underlying uniformly locally finite metric spaces. In [2], the authors provided an affirmative answer to the rigidity problem in the case of Morita equivalences.
@@
    One shows that the unitaries $U$ and $U^{*}$ implementing $\Phi$ are coarse-like [^5] In [15, 14], this notion is also referred to as weakly approximately controlled. (see [7, Theorem 3.5]). Recall from [7, Definition 3.1] that an operator
@@
    One shows that, for every $\delta\in(0,1)$, there exist constants $R,r\geq 0$ such that the subset
    $\quad\qquad f_{\delta,R,r}^{U}=\bigcup\left\{B\times A\;\middle|\;\operatorname{diam}(B)\leq R,\ \operatorname{diam}(A)\leq r,\ \bigl\|\mathbbm{1}_{B}\,U\,\mathbbm{1}_{A}\bigr\|>\delta\right\}$
@@ 1.3. Uniformly finite homology
Theorem 1.14 ([28, Corollary 3.1 and Corollary 3.2]).
@@
In particular, if $(X,d)$ is an amenable, coarsely connected, uniformly locally finite metric space, then $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is a non-zero real vector space; see [28, Theorem D]. It also follows from the above result that the 0th uniformly finite homology group is always torsion-free.
@@
The map $\alpha_{0}$ coincides with the 0th comparison map appearing in Conjecture 0.2 for coarse groupoids. It remains an open problem whether this map is split injective in general.
@@ 2. Uniform covering isometries
Let $H$ be a separable Hilbert space. It is well known (for instance, see [29, Section 4.3]) that, for a proper coarse map between two uniformly locally finite metric spaces $f\colon(X,d)\to(Y,\partial)$, there exists an isometry $S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$, called a covering isometry of $f$, such that the support of $S$ is asymptotic to the graph of $f$ and $\operatorname{Ad}_{S}$ restricts to a *-homomorphism between the associated Roe algebras. Moreover, the induced map in $K$-theory depends only on $f$. However, covering isometries do not, in general, induce *-homomorphisms between uniform Roe algebras or their stabilisations. In this section, we introduce a modification of the above notion adapted to the setting of stabilised uniform Roe algebras.
@@
Since $C^{*}_{u}(X)\otimes\mathbb{K}$ is a non-degenerate $C^{*}$-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$, it follows from [4, II 7.3.5] that its multiplier algebra can be identified with the $C^{*}$-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$ given by
@@
For the case of coarse equivalences, the existence of covering unitaries was established in [8]. The following lemma is a refinement of [8, Theorem 4], it establishes the existence of a uniform cover for any uniformly finite-to-one coarse map.
@@
Moreover, by [6, Theorem G], there is a natural isomorphism between the homology of coarse groupoids and uniformly finite homology. As the uniform Roe algebra is naturally isomorphic to the reduced groupoid $C^{*}$-algebra of the associated coarse groupoid, the commutativity of the diagram in Lemma 2.5 follows.
@@ 3. Rigidity of uniform Roe algebras versus injectivity of $\alpha_{0}$
In this section, we establish a close connection between the injectivity of the 0th comparison map for coarse groupoids and the bijective rigidity problem. In contrast to arguments involving a single space, our approach necessarily requires the consideration of subspaces and doublings. The following definition formalises the framework needed for this purpose.
@@
A collection $\mathcal{C}$ of uniformly locally finite metric spaces is called admissible if it is closed under taking arbitrary doublings and subspaces. An admissible collection $\mathcal{C}$ is said to satisfy bijective rigidity if, for any uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ and any isomorphism $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$, the induced coarse equivalence $f_{\Phi}\colon(X,d)\to(Y,\partial)$ provided by Theorem 1.7 is close to a bijective coarse equivalence.
For example, let $\mathcal{C}(\mathrm{A})$ denote the collection of all uniformly locally finite metric spaces that satisfy Property A, and let $\mathcal{C}(\mathrm{CE})$ denote the collection of all uniformly locally finite metric spaces that admit a coarse embedding into a Hilbert space. Since both properties are preserved under taking subspaces and under passage to coarsely equivalent spaces, the collections $\mathcal{C}(\mathrm{A})$ and $\mathcal{C}(\mathrm{CE})$ are admissible. Clearly, the collection of all uniformly locally finite metric spaces is also admissible. The following theorem is the main result of this section.
Theorem A.
Let $\mathcal{C}$ be an admissible collection of uniformly locally finite metric spaces. Then $\mathcal{C}$ satisfies bijective rigidity if and only if, for every uniformly locally finite metric space $(X,d)$ in $\mathcal{C}$, the 0th comparison map
$\alpha_{0}\colon H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\to K_{0}(C^{*}_{u}(X))$
is injective.
Note that, for ample groupoids, the injectivity of $\alpha_{0}$ is invariant under Morita equivalence; hence, the injectivity of $\alpha_{0}$ is preserved by passing to coarsely equivalent spaces. In particular, the 0th comparison map for $(X,d)$ is injective if and only if it is injective for any (some) doubling of $(X,d)$. On the other hand, it is not evident whether injectivity passes to subspaces. Before proceeding with the proof of Theorem A, we need to establish a connection between the $K$-theory of an isomorphism $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$, and the $K$-theory of the induced coarse equivalence $f_{\Phi}$.
@@
In the proof of Proposition 3.2, we showed that there exists a unitary $\tilde{U}$ in the multiplier algebra of $C^{*}_{u}(X)\otimes\mathbb{K}$ which intertwines the *-isomorphisms $\Phi\otimes\operatorname{id}_{H}$ and $\operatorname{Ad}_{U}$, where $U$ is a uniform cover of $f_{\Phi}$. Recall from [17, Theorem 2.5] that for any unital $C^{*}$-algebra $A$, the unitary group of the multiplier algebra of $A\otimes\mathbb{K}$ is contractible. Consequently, there exists a continuous path of unitaries $u_{t}\in\mathcal{M}(C^{*}_{u}(X)\otimes\mathbb{K})$ such that $u_{0}=\tilde{U}$ and $u_{1}=1$. In particular, the *-isomorphisms $\Phi\otimes\operatorname{id}_{H}$ and $\operatorname{Ad}_{U}$ are homotopic.
@@
and the right-hand side is exactly the $K_{0}$-class of $\mathbbm{1}_{Y}$. Since $\alpha_{0}^{Y}$ is injective, it follows that $H_{0}^{\text{uf}}(f_{\Phi})$ preserves the fundamental class. By Theorem 1.12, there exists a bijective coarse equivalence that is close to $f_{\Phi}$. Consequently, the collection $\mathcal{C}$ satisfies bijective rigidity.
Conversely, let $\mathcal{C}$ be an admissible collection satisfying bijective rigidity, and let $(X,d)\in\mathcal{C}$. Suppose that $\beta\in\ker(\alpha_{0}^{X})$. By passing, if necessary, to a doubling of $(X,d)$, for some $n\geq 1$ the class $\beta$ is given by
@@
Since $\mathcal{C}$ satisfies bijective rigidity, this coarse equivalence may be chosen to be bijective. Moreover, because the partial isometry $s$ is approximable, the associated coarse equivalence $f_{\Phi}$ is close to the identity map on $X$. In particular, $f_{\Phi}$ is a partial translation. Hence the indicator functions of $A\sqcup X^{(n+\ell)}$ and $B\sqcup X^{(n+\ell)}$ define the same class in $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$. By substracting the homology class of the indicator function of $X^{(n+\ell)}$, we conclude that $\mathbbm{1}_{A}]=[\mathbbm{1}_{B}]$, and therefore $\beta=0$. This shows that $\alpha_{0}^{X}$ is injective. ∎
@@
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ be an isomorphism. Denote by $f_{\Phi}$ the coarse equivalence provided by Theorem 1.7. Assume that the 0th comparison map $\alpha_{0}^{X}$ is injective. By Lemma 2.5, since $H_{0}^{\mathrm{uf}}(f_{\Phi})$ and $K_{0}(f_{\Phi})$ are isomorphisms, it follows that the 0th comparison map $\alpha_{0}^{Y}$ is also injective. Consequently, by the proof of Theorem A, the coarse equivalence $f_{\Phi}$ is close to a bijective coarse equivalence. ∎
@@
It follows that injectivity of $\alpha_{0}^{X}\otimes\operatorname{id}_{\mathbb{Q}}$ implies injectivity of $\alpha_{0}^{X}$. Consequently, in Theorem A and its corollaries, it suffices to assume rational injectivity of the 0th comparison map.
@@ 4. Consequences
In [27, Theorem E] it is shown that the class $\mathcal{C}(\mathrm{A})$ of uniformly locally finite metric spaces satisfying property A (equivalently, the class of amenable coarse groupoids) satisfies bijective rigidity. By applying Theorem A, we obtain that the 0th comparison map is injective for all uniformly locally finite spaces that satisfy Property A. The authors of [6] (see Corollary C) proved that for any ample, principal, second-countable, locally compact, Hausdorff groupoid $\mathcal{G}$ with dynamic asymptotic dimension at most $2$, such that $H_{2}(\mathcal{G};\mathbb{Z})$ is finitely generated, there are isomorphisms
@@
They also demonstrated that the result applies to coarse groupoids. For coarse groupoids, the dynamical asymptotic dimension coincides with the asymptotic dimension of the underlying metric space. Consequently, the above isomorphisms hold for all uniformly locally finite metric spaces of asymptotic dimension at most $2$ with finitely generated $H_{2}^{\text{uf}}(X;\mathbb{Z})$. Since every uniformly locally finite metric space of finite asymptotic dimension has property A, our results may be viewed as a partial generalisation of the results of [6]. Recently, Vignati showed in [26] that bijective rigidity holds for all uniformly locally finite metric spaces; equivalently, the collection of all uniformly locally finite metric spaces satisfies bijective rigidity.
Corollary C.
@@
By [26], the collection of all uniformly locally finite metric spaces satisfies bijective rigidity. It therefore follows from Theorem A that, for any uniformly locally finite metric space $(X,d)$, the comparison map $\alpha_{0}$ is injective. Suppose now that $(X,d)$ is coarsely connected. By Theorem 1.14, the canonical map
@@
Let $G$ be a finitely generated group equipped with the word-length metric. By [11, Proposition A.10], its 0th uniformly finite homology group canonically identifies with the group homology of $G$ with coefficients in $\ell^{\infty}(G,\mathbb{R})$. Since Cayley graphs of finitely generated groups are coarsely connected, we obtain the following application.
Corollary D.
@@
Corollary C relies on the fact that real vector spaces are injective objects in the category of abelian groups. In general, however, $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ need not be injective. For example, if $X$ is finite, then $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is isomorphic to $\mathbb{Z}$. As another example, consider the subset $\mathbb{N}_{s}:=\{n^{2}\mid n\in\mathbb{N}\}$ of $\mathbb{N}$ equipped with the induced metric. This space is sometimes referred to as the sparse natural numbers. One readily verifies that
$H_{0}^{\mathrm{uf}}(\mathbb{N}_{s};\mathbb{Z})\cong\ell^{\infty}(\mathbb{N},\mathbb{Z})/c_{0}(\mathbb{N},\mathbb{Z}).$
The latter group is not divisible and therefore is not injective. Note that, for any abelian group $A$, the group $A\otimes_{\mathbb{Z}}\mathbb{Q}$ is a $\mathbb{Q}$-vector space and therefore injective. It is an immediate consequence of Corollary C and Remark 3.4 that the rational comparison map
@@
where the left-hand side denotes the topological $K$-theory of $\mathcal{G}$. In particular, assuming the rational injectivity of the Baum–Connes assembly map, one recovers the rational split-injectivity of the comparison map. As coarse groupoids are, in general, not second-countable, one has to provide an additional limiting argument akin to the one used in [6, Section 5.3] to apply the machinery of [18]. This might serve as an alternative proof of the split-injectivity of the rational comparison map for coarse groupoids whose Baum–Connes assembly map is rationally injective. For example, by the results of [23, 25], the class $\mathcal{C}(\mathrm{CE})$ of uniformly locally finite metric spaces admitting a coarse embedding into a Hilbert space satisfies the Baum–Connes conjecture with coefficients; in particular, for these spaces, the Baum-Connes assembly map is rationally injective.
@@
such that the 0th comparison map $\alpha_{0}$ factors through $\alpha^{0}$ and the Baum-Connes map. In the case of coarse groupoids, Corollary C establishes an unconditional low-dimensional injectivity of the Baum-Connes map. We shall return to this perspective in our future work.
Acknowledgements
The author is grateful to his advisors, Alessandro Vignati and Romain Tessera, for their advice on the presentation of this work. He also thanks Ján Špakula and the University of Southampton for their hospitality during his research visit. The author is indebted to Rufus Willett for valuable comments on Remark 4.1. This work was partially supported by the ECOST Actions CaLISTA travel grant.

### v2

@@ 1.2. Uniform Roe algebras and rigidity
The rigidity problem for uniform Roe algebras asks whether Morita equivalences (respectively, isomorphisms) of uniform Roe algebras induce coarse equivalences (respectively, bijective coarse equivalences) between the underlying uniformly locally finite metric spaces. In [2], the authors provided an affirmative answer to the rigidity problem in the case of Morita equivalences.
@@
    One shows that the unitaries $U$ and $U^{*}$ implementing $\Phi$ are coarse-like [^7] In [16, 15], this notion is also referred to as weakly approximately controlled. (see [6, Theorem 3.5]). Recall from [6, Definition 3.1] that an operator
@@
    One shows that, for every $\delta\in(0,1)$, there exist constants $R,S\geq 0$ such that the subset
    $\quad\qquad f_{\delta,R,S}^{U}=\bigcup\left\{B\times A\;\middle|\;\operatorname{diam}(B)\leq R,\ \operatorname{diam}(A)\leq S,\ \bigl\|\mathbbm{1}_{B}\,U\,\mathbbm{1}_{A}\bigr\|>\delta\right\}$
@@ 1.3. Uniformly finite homology
Theorem 1.14 ([29, Corollary 3.1 and Corollary 3.2]).
@@
In particular, if $(X,d)$ is an amenable, coarsely connected, uniformly locally finite metric space, then $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is a non-zero real vector space; see [29, Theorem D]. It also follows from the above result that the 0th uniformly finite homology group is always torsion-free.
@@
The map $\alpha_{0}$ coincides with the 0th comparison map appearing in Conjecture Conjecture for coarse groupoids. It remains an open problem whether this map is split injective in general.
@@ 2. Uniform covering isometries
Let $H$ be a separable Hilbert space. It is well known (for instance, see [30, Section 4.3]) that, for a proper coarse map between two uniformly locally finite metric spaces $f\colon(X,d)\to(Y,\partial)$, there exists an isometry $S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$, called a covering isometry of $f$, such that the support of $S$ is asymptotic to the graph of $f$ and $\operatorname{Ad}_{S}$ restricts to a *-homomorphism between the associated Roe algebras. Moreover, the induced map in $K$-theory depends only on $f$. However, covering isometries do not, in general, induce *-homomorphisms between uniform Roe algebras or their stabilisations. In this section, we introduce a modification of the above notion adapted to the setting of stabilised uniform Roe algebras.
@@
Since $C^{*}_{u}(X)\otimes\mathbb{K}$ is a non-degenerate $C^{*}$-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$, it follows from [3, II 7.3.5] that its multiplier algebra can be identified with the $C^{*}$-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$ given by
@@
For the case of coarse equivalences, the existence of covering unitaries was established in [7]. The following lemma is a refinement of [7, Theorem 4], it establishes the existence of a uniform cover for any uniformly finite-to-one coarse map.
@@
Moreover, by [5, Theorem G], there is a natural isomorphism between the homology of coarse groupoids and uniformly finite homology. As the uniform Roe algebra is naturally isomorphic to the reduced groupoid $C^{*}$-algebra of the associated coarse groupoid, the commutativity of the diagram in Lemma 2.5 follows.
@@ 3. Rigidity of uniform Roe algebras versus injectivity of $\alpha_{0}$
In this section, we establish a close connection between the injectivity of the 0th comparison map for coarse groupoids and the bijective rigidity problem. In contrast to arguments involving a single space, our approach necessarily requires the consideration of subspaces and doublings.
@@
A collection $\mathcal{C}$ of uniformly locally finite metric spaces is called admissible if it is closed under taking arbitrary doublings and subspaces.
For example, let $\mathcal{C}(\mathrm{A})$ denote the collection of all uniformly locally finite metric spaces that satisfy Property A, and let $\mathcal{C}(\mathrm{CE})$ denote the collection of all uniformly locally finite metric spaces that admit a coarse embedding into a Hilbert space. Since both properties are preserved under taking subspaces and under passage to coarsely equivalent spaces, the collections $\mathcal{C}(\mathrm{A})$ and $\mathcal{C}(\mathrm{CE})$ are admissible. Clearly, the collection of all uniformly locally finite metric spaces is also admissible. The following lemma is the main result of this section.
Lemma 3.2 (Lemma A).
  1. (1)
    For every uniformly locally finite metric space $(X,d)$ in $\mathcal{C}$, the 0th comparison map $\alpha_{0}\colon H^{\mathrm{uf}}_{0}(X;\mathbb{Z})\to K_{0}\bigl(C^{*}_{u}(X)\bigr)$ is injective;
  2. (2)
    For any uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ such that $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 1.7 is close to a bijective coarse equivalence.
Note that, for ample groupoids, the injectivity of $\alpha_{0}$ is invariant under Morita equivalence; hence, the injectivity of $\alpha_{0}$ is preserved by passing to coarsely equivalent spaces. In particular, the 0th comparison map for $(X,d)$ is injective if and only if it is injective for any (some) doubling of $(X,d)$. On the other hand, it is not evident whether injectivity passes to subspaces. Before proceeding with the proof of Lemma 3.2, we need to establish a connection between the $K$-theory of an isomorphism $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$, and the $K$-theory of the coarse equivalence $f_{\Phi}$ induced by Theorem 1.7.
@@
In the proof of Proposition 3.3, we showed that there exists a unitary $\tilde{U}$ in the multiplier algebra of $C^{*}_{u}(X)\otimes\mathbb{K}$ which intertwines the *-isomorphisms $\Phi\otimes\operatorname{id}_{H}$ and $\operatorname{Ad}_{U}$, where $U$ is a uniform cover of $f_{\Phi}$. Recall from [18, Theorem 2.5] that for any unital $C^{*}$-algebra $A$, the unitary group of the multiplier algebra of $A\otimes\mathbb{K}$ is contractible. Consequently, there exists a continuous path of unitaries $u_{t}\in\mathcal{M}(C^{*}_{u}(X)\otimes\mathbb{K})$ such that $u_{0}=\tilde{U}$ and $u_{1}=1$. In particular, the *-isomorphisms $\Phi\otimes\operatorname{id}_{H}$ and $\operatorname{Ad}_{U}$ are homotopic.
@@
and the right-hand side is exactly the $K_{0}$-class of $\mathbbm{1}_{Y}$. Since $\alpha_{0}^{Y}$ is injective, it follows that $H_{0}^{\text{uf}}(f_{\Phi})$ preserves the fundamental class. By Theorem 1.12, there exists a bijective coarse equivalence that is close to $f_{\Phi}$. Consequently, the coarse equivalence given by Theorem 1.7 is close to a bijective coarse equivalence.
Conversely, let $\mathcal{C}$ be an admissible collection of uniformly locally finite metric spaces such that for any $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ satisfing $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 1.7 is close to a bijective coarse equivalence. Let $(X,d)\in\mathcal{C}$, and let $\beta\in\ker(\alpha_{0}^{X})$. By passing, if necessary, to a doubling of $(X,d)$, for some $n\geq 1$ the class $\beta$ is given by
@@
Recall that, up to closeness, for some (equivalently, any) $\delta\in(0,1)$, there exist constants $R,P\geq 0$ such that the relation $f_{\Phi}$ has the following form:
$f_{\Phi}=\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|>\delta\}.$
Since $s$ is approximable, for every $\varepsilon>0$ there exists a controlled propagation operator $t$ such that $\|s-t\|<\varepsilon$. Choose $\varepsilon>0$ with $\varepsilon<\delta$. If $C,D\subset X$ satisfy $\|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|>\delta$, then
$\delta<\|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|\leq\|\mathbbm{1}_{C}(s-t)\mathbbm{1}_{D}\|+\|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|<\varepsilon+\|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|.$
In particular, $\|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|>\delta-\varepsilon$. Consequently, we obtain a chain of inclusions
$\begin{split}f_{\Phi}=\;&\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|>\delta\}\\ \subset\;&\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|>\delta-\varepsilon\}\\ \subset\;&\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|\neq 0\}\\ \subset\;&E_{R}\circ\operatorname{supp}(t)\circ E_{P}.\end{split}$
Since $t$ has controlled propagation, its support is an entourage. As entourages are closed under composition, the latter set is also an entourage. In particular, the relation $f_{\Phi}$ is contained in an entourage; therefore, it is close to the identity map on $X$. By assumption, $f_{\Phi}$ can be chosen to be bijective; in particular, $f_{\Phi}$ is a partial translation. Hence the indicator functions of $A\sqcup X^{(n+\ell)}$ and $B\sqcup X^{(n+\ell)}$ define the same class in $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$. By substracting the homology class of the indicator function of $X^{(n+\ell)}$, we conclude that $\mathbbm{1}_{A}]=[\mathbbm{1}_{B}]$, and therefore $\beta=0$. This shows that $\alpha_{0}^{X}$ is injective. ∎
@@
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ be an isomorphism. Denote by $f_{\Phi}$ the coarse equivalence provided by Theorem 1.7. Assume that the 0th comparison map $\alpha_{0}^{X}$ is injective. By Lemma 2.5, since $H_{0}^{\mathrm{uf}}(f_{\Phi})$ and $K_{0}(f_{\Phi})$ are isomorphisms, it follows that the 0th comparison map $\alpha_{0}^{Y}$ is also injective. Consequently, by the proof of Lemma 3.2, the coarse equivalence $f_{\Phi}$ is close to a bijective coarse equivalence. ∎
@@
It follows that injectivity of $\alpha_{0}^{X}\otimes\operatorname{id}_{\mathbb{Q}}$ implies injectivity of $\alpha_{0}^{X}$. Consequently, in Lemma 3.2 and its corollaries, it suffices to assume rational injectivity of the 0th comparison map.
@@ 4. Consequences
In [28, Theorem E] it is shown that the class $\mathcal{C}(\mathrm{A})$ of uniformly locally finite metric spaces satisfying property A (equivalently, the class of amenable coarse groupoids) satisfies bijective rigidity. By applying Lemma 3.2, we obtain that the 0th comparison map is injective for all uniformly locally finite spaces that satisfy Property A. The authors of [5] (see Corollary C) proved that for any ample, principal, second-countable, locally compact, Hausdorff groupoid $\mathcal{G}$ with dynamic asymptotic dimension at most $2$, such that $H_{2}(\mathcal{G};\mathbb{Z})$ is finitely generated, there are isomorphisms
@@
They also demonstrated that the result applies to coarse groupoids. For coarse groupoids, the dynamical asymptotic dimension coincides with the asymptotic dimension of the underlying metric space. Consequently, the above isomorphisms hold for all uniformly locally finite metric spaces of asymptotic dimension at most $2$ with finitely generated $H_{2}^{\text{uf}}(X;\mathbb{Z})$. Since every uniformly locally finite metric space of finite asymptotic dimension has property A, our results may be viewed as a partial generalisation of the results of [5]. Recently, independently of this work, Vignati showed in [27] that bijective rigidity holds for all uniformly locally finite metric spaces (see Theorem 0.2). Moreover, the bijective coarse equivalence constructed therein is close to the coarse equivalence provided by Theorem 1.7. We obtain the following theorem.
Theorem 4.1 (Theorem B).
@@
By [27], the second assertion of Lemma 3.2 holds; therefore, for any uniformly locally finite metric space $(X,d)$, the comparison map $\alpha_{0}$ is injective. Suppose now that $(X,d)$ is coarsely connected. By Theorem 1.14, the canonical map
@@
Let $G$ be a finitely generated group equipped with the word-length metric. By [10, Proposition A.10], its 0th uniformly finite homology group canonically identifies with the group homology of $G$ with coefficients in $\ell^{\infty}(G,\mathbb{R})$. Since Cayley graphs of finitely generated groups are coarsely connected, we obtain the following application.
Corollary 4.2 (Corollary C).
@@
Corollary 4.2 relies on the fact that real vector spaces are injective objects in the category of abelian groups. In general, however, $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ need not be injective. For example, if $X$ is finite, then $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is isomorphic to $\mathbb{Z}$. As another example, consider the subset $\mathbb{N}_{s}:=\{n^{2}\mid n\in\mathbb{N}\}$ of $\mathbb{N}$ equipped with the induced metric. This space is sometimes referred to as the sparse natural numbers. One readily verifies that
$H_{0}^{\mathrm{uf}}(\mathbb{N}_{s};\mathbb{Z})\cong\ell^{\infty}(\mathbb{N},\mathbb{Z})/\{f\in c_{0}(\mathbb{N},\mathbb{Z})\mid\sum_{n\in\mathbb{N}}f(n)=0\}.$
The latter group is not divisible and therefore is not injective. Note that, for any abelian group $A$, the group $A\otimes_{\mathbb{Z}}\mathbb{Q}$ is a $\mathbb{Q}$-vector space and therefore injective. It is an immediate consequence of Theorem 4.1 and Remark 3.6 that the rational comparison map
@@
where the left-hand side denotes the topological $K$-theory of $\mathcal{G}$. In particular, assuming the rational injectivity of the Baum–Connes assembly map, one recovers the rational split-injectivity of the comparison map. As coarse groupoids are, in general, not second-countable, one has to provide an additional limiting argument akin to the one used in [5, Section 5.3] to apply the machinery of [19]. This might serve as an alternative proof of the split-injectivity of the rational comparison map for coarse groupoids whose Baum–Connes assembly map is rationally injective. For example, by the results of [23, 26], the class $\mathcal{C}(\mathrm{CE})$ of uniformly locally finite metric spaces admitting a coarse embedding into a Hilbert space satisfies the Baum–Connes conjecture with coefficients; in particular, for these spaces, the Baum-Connes assembly map is rationally injective.
@@
such that the 0th comparison map $\alpha_{0}$ factors through $\alpha^{0}$ and the Baum-Connes map. In the case of coarse groupoids, Theorem 4.1 establishes an unconditional low-dimensional injectivity of the Baum-Connes map. We shall return to this perspective in our future work.